Birch Company normally produces and sells 45,000 units of RG-6 each month. The selling price is $30 per unit, variable costs are $10 per unit, fixed manufacturing overhead costs total $190,000 per month, and fixed selling costs total $40,000 per month.

Employment-contract strikes in the companies that purchase the bulk of the RG-6 units have caused Birch Company’s sales to temporarily drop to only 10,000 units per month. Birch Company estimates that the strikes will last for two months, after which time sales of RG-6 should return to normal. Due to the current low level of sales, Birch Company is thinking about closing down its own plant during the strike, which would reduce its fixed manufacturing overhead costs by $48,000 per month and its fixed selling costs by 10%. Start-up costs at the end of the shutdown period would total $15,000. Because Birch Company uses Lean Production methods, no inventories are on hand.

Required: 1. What is the financial advantage (disadvantage) if Birch closes its own plant for two months? 2. Should Birch close the plant for two months? 3. At what level of unit sales for the two-month period would Birch Company be indifferent between closing the plant or keeping it open?
Answer #1

Solution:

1) We analyse two situations here to find whether there is financial advantage or disadvantage if Birch closes its own plant for two months.

1 st situation : Birch company does not shut down and continues to operate.

Sales (10,000 units x 2 months x 30) $ 600,000
Less: Variable costs ( 10,000 units x 2 months x 10) $ 200,000
Total contribution $ 400,000
Less: Fixed manufacturing overhead ( $190,000 x 2) $ 380,000
Less: Fixed Selling costs ( $ 40,000 x 2) $ 80,000
Net Income/ (Loss) ($ 60,000)

2nd situation : Birch company decided to shut down its plant for 2 months during strike.

Unavoidable Fixed Manufacturing overhead (190,000 - 48,000) x 2)) ($ 284,000)
Unavoidable Fixed selling costs (40,000 x 90 % x 2) ($ 72,000)
Start-up costs at the end of the shutdown period ($15,000)
Total loss due to shut down ($ 371,000)

Since the loss is greater when the company shut downs it plant it faces a financial disadvantage.

Financial Disadvantage = 371,000 - 60,000 = $ 311,000

b) No, Birch company should not close its plant but continue to operate as the unavoidable expenses during shutdown would be greater than the losses from operation.

c) Birch company would be indifferent between closing the plant or keeping it open when the losses from operation and shutdown are equal. So, from back calculations we find out at what level of unit sales loss from operation is $ 371,000.

If the loss from operation is $ 371,000 , we calculate contribution by removing the effect of fixed manufacturing overhead costs and fixed selling costs from Net loss.

Contribution would be = 371,000 - 80,000 - 380,000 = $ 89,000

Contribution per unit is = selling price per unit - variable cost per unit

= 30 - 10 = 20 per unit

Number of units required = Total contribution / contribution per unit

= $ 89,000 / 20

= 4450 units i.e, 4450/2 = 2225 units per month.